OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x2 = 20y

OpenStudy (anonymous):

it is in standard form x2=4by hence hre b=20/4=5 thus verex (0,0) focus(0,5) directrix is y+5=0 and focal width=4*5 units=20 units

OpenStudy (anonymous):

Vertex: (0, 0); Focus: (0, 5); Directrix: y = -5; Focal width: 20 Vertex: (0, 0); Focus: (5, 0); Directrix: x = 5; Focal width: 5 Vertex: (0, 0); Focus: (5, 0); Directrix: y = 5; Focal width: 80 Vertex: (0, 0); Focus: (0, -5); Directrix: x = -5; Focal width: 80

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

Vertex: (0, 0); Focus: (0, 5); Directrix: y = -5; Focal width: 20

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.

OpenStudy (anonymous):

y =1/28 x2 x = 1/28 y2 -28y = x2 y2 = 14x

OpenStudy (anonymous):

@matricked